# Learning to detect continuous gravitational waves: an open data-analysis competition

### v1

@@ Acknowledgements
We thank Rafel Jaume, David Keitel, Narenraju Nagarajan, Alicia M. Sintes, and Daniel Williams for support and discussions, and Christopher Berry, Lorenzo Mirasola, and Karl Wette for comments on the manuscript. We thank Kaggle/Google for contributing to the competition prize, and acknowledge support from COST Action CA17137 (G2Net). RT thanks Graham Woan and the Institute for Gravitational Research at the University of Glasgow for their hospitality during the development of this work. RT is supported by ERC Starting Grant No. 945155–GWmining; Cariplo Foundation Grant No. 2021-0555; MUR PRIN Grant No. 2022-Z9X4XS; MUR Grant “Progetto Dipartimenti di Eccellenza 2023-2027” (BiCoQ); Italian-French University (UIF/UFI) Grant No. 2025-C3-386; the ICSC National Research Centre funded by NextGenerationEU; the Universitat de les Illes Balears (UIB); the Spanish Agencia Estatal de Investigación grants PID2022-138626NB-I00, RED2024-153978-E, RED2024-153735-E, funded by MICIU/AEI/10.13039/501100011033 and the ERDF/EU; and the Comunitat Autònoma de les Illes Balears through the Conselleria d’Educació i Universitats with funds from the European Union - NextGenerationEU/PRTR-C17.I1 (SINCO2022/6719) and from the European Union - European Regional Development Fund (ERDF) (SINCO2022/18146). MJW is supported by the Science and Technology Facilities Council [2285031, ST/X002225/1, ST/Y004876/1], and the University of Portsmouth. JB and CM are supported by STFC grant ST/V005634/1. Computational work was performed at MareNostrum5 with technical support provided by the Barcelona Supercomputing Center (RES-FI-2024-3-0013, RES-FI-2025-1-0022), and at CINECA with allocations through INFN and Bicocca. This research has made use of data or software obtained from the Gravitational Wave Open Science Center (gw-openscience.org), a service of LIGO Laboratory, the LIGO Scientific Collaboration, the Virgo Collaboration, and KAGRA. This paper has been assigned document number LIGO-P2200295.
@@ Appendix A First Kaggle Competition
The dataset comprised simulated measurements from the three-detector network of advanced interferometers: LIGO-Hanford, LIGO-Livingston, and Virgo. Each training example consisted of three time series, corresponding to the three detectors, covering two seconds sampled at 2048 Hz. Some samples contained only Gaussian noise, while others were augmented with simulated GW signals. The simulated BBH signals were generated using IMRPhenomPv2 [127] and varied across 15 astrophysically relevant parameters, including component masses, spins, binary orientation, sky location, distance, polarization, merger phase, and arrival time. Although the distributions of these parameters reflected realistic astrophysical populations, the overall signal amplitudes were adjusted to produce a mix of challenging and easier detection cases. Most signals were designed to lie near the typical sensitivity threshold of matched-filter searches, around an SNR of 8.
The challenge attracted broad participation, with 1,501 individuals forming 1,219 teams. Competitors applied a variety of machine learning approaches, from convolutional and recurrent neural networks to hybrid architectures, often incorporating data augmentation and ensemble methods to improve generalization. The competition validated claims in the literature [36, 128, 129] that deep learning methods could achieve competitive detection performance on realistically simulated multi-detector GW datasets. Moreover, the challenge provided a benchmark for future efforts in applying modern machine learning techniques to gravitational-wave detection, highlighting both the potential and limitations of data-driven approaches in this domain.

### v2

@@ Acknowledgements
We are grateful to the all Kaggle competitors for their invaluable input and engagement throughout the competition. We thank Rafel Jaume, David Keitel, Narenraju Nagarajan, Alicia M. Sintes, and Daniel Williams for support and discussions, and Christopher Berry, Lorenzo Mirasola, and Karl Wette for comments on the manuscript. We thank Kaggle/Google for contributing to the competition prize, and acknowledge support from COST Action CA17137 (G2Net). RT thanks Graham Woan and the Institute for Gravitational Research at the University of Glasgow for their hospitality during the development of this work. RT is supported by ERC Starting Grant No. 945155–GWmining; Cariplo Foundation Grant No. 2021-0555; MUR PRIN Grant No. 2022-Z9X4XS; MUR Grant “Progetto Dipartimenti di Eccellenza 2023-2027” (BiCoQ); Italian-French University (UIF/UFI) Grant No. 2025-C3-386; the ICSC National Research Centre funded by NextGenerationEU; the Universitat de les Illes Balears (UIB); the Spanish Agencia Estatal de Investigación grants PID2022-138626NB-I00, RED2024-153978-E, RED2024-153735-E, funded by MICIU/AEI/10.13039/501100011033 and the ERDF/EU; and the Comunitat Autònoma de les Illes Balears through the Conselleria d’Educació i Universitats with funds from the European Union - NextGenerationEU/PRTR-C17.I1 (SINCO2022/6719) and from the European Union - European Regional Development Fund (ERDF) (SINCO2022/18146). MJW is supported by the Science and Technology Facilities Council [2285031, ST/X002225/1, ST/Y004876/1], and the University of Portsmouth. JB and CM are supported by STFC grant ST/V005634/1. Computational work was performed at MareNostrum5 with technical support provided by the Barcelona Supercomputing Center (RES-FI-2024-3-0013, RES-FI-2025-1-0022), at CINECA with allocations through INFN and Bicocca. This research has made use of data or software obtained from the Gravitational Wave Open Science Center (gw-openscience.org), a service of LIGO Laboratory, the LIGO Scientific Collaboration, the Virgo Collaboration, and KAGRA. This paper has been assigned document number LIGO-P2200295.
@@ Appendix A First Kaggle Competition
The dataset comprised simulated measurements from the three-detector network of advanced interferometers: LIGO-Hanford, LIGO-Livingston, and Virgo. Each training example consisted of three time series, corresponding to the three detectors, covering two seconds sampled at 2048 Hz. Some samples contained only Gaussian noise, while others were augmented with simulated GW signals. The simulated BBH signals were generated using IMRPhenomPv2 [Hannam:2013oca] and varied across 15 astrophysically relevant parameters, including component masses, spins, binary orientation, sky location, distance, polarization, merger phase, and arrival time. Although the distributions of these parameters reflected realistic astrophysical populations, the overall signal amplitudes were adjusted to produce a mix of challenging and easier detection cases. Most signals were designed to lie near the typical sensitivity threshold of matched-filter searches, around an SNR of 8.
The challenge attracted broad participation, with 1,501 individuals forming 1,219 teams. Competitors applied a variety of machine learning approaches, from convolutional and recurrent neural networks to hybrid architectures, often incorporating data augmentation and ensemble methods to improve generalization. The competition validated claims in the literature [Gabbard:2017lja, 2018PhRvD..97d4039G, 2019PhRvD.100f3015G] that deep learning methods could achieve competitive detection performance on realistically simulated multi-detector GW datasets. Moreover, the challenge provided a benchmark for future efforts in applying modern machine learning techniques to gravitational-wave detection, highlighting both the potential and limitations of data-driven approaches in this domain.